# Fair Play in the Newsroom: Actor-Based Filtering Gender Discrimination in Text Corpora

### v1

@@ 5 Corpus-Balancing of taz2024full
We use the taz2024full corpus Urchs et al. (2025), comprising over 1.8 million articles from the German left-leaning newspaper taz (1980–2024), previously used to analyse gender representation. In the unfiltered taz2024full corpus, we detect female- and male-coded actors in 1,834,018 articles. Gender representation is clearly imbalanced: Men dominate across both mention frequency and actor counts (Figure 1). These differences are not only quantitative but also reflected in patterns of agency and authority.
Figure 1: Percentage of male- and female-coded references over time before filtering. Fluctuations in the early years are due to the small number of articles available from the 1980s, which can lead to disproportionate weight for single-gender articles.
Figure 2 shows that male actors are more frequently quoted directly, while female actors appear more often in indirect quotes. Direct quotations often attribute more authority to the speaker and allow them to appear in active, public-facing roles. Indirect quotes, by contrast, reduce visibility and typically appear in paraphrased or backgrounded contexts.
Figure 2: Distribution of quotation styles by gender before filtering. Early-year fluctuations are again attributable to low article counts.
Figure 3 further underlines this pattern. Men occur more frequently in subject positions, while women appear comparatively more often as objects. Subject roles in grammar typically denote agency and narrative control, whereas object positions signal reduced agency and passivity within the sentence structure.
Figure 3: Distribution of syntactic roles by gender before filtering. Early-year fluctuations are again attributable to low article counts.
Sentiment analysis reveals a consistent gap: The average sentiment towards female-coded actors is more negative in almost all years (Figure 4). While sentiment values are generally close to neutral, the persistent divergence reinforces the broader imbalance.
Figure 4: Average sentiment associated with male- and female-coded actors before filtering.
The early-year spikes in Figures 1, 2, and 3 are artefacts of low data density. In years with very few articles or actor mentions, a single gender may dominate, resulting in sharp fluctuations that do not reflect structural bias but instead data sparsity. We retain these years for completeness, but advise caution when interpreting early data points and refrain from doing this ourselves.
Finally, the overall distribution of gender representation per article (Figure 5) is highly polarised: many articles reference either only male-coded or only female-coded actors. This reinforces the need for corpus-level balancing, as it shows that article-level imbalance is not just a matter of aggregate statistics but of individual article composition.
@@
During the first text-level filtering step, we exclude 279,772 articles based on four framing asymmetries: sentiment gap, quote imbalance, subject/object ratio, and representation. We use the default values described in Section 4.2. The updated diagnostic view shows reduced polarisation, and the gender ratio distribution shifts closer to balance (Figure 6).
Figure 6: Distribution of gender ratios across articles after text-level filtering.
Subsequently, we apply corpus-level balancing, excluding an additional 17,815 articles to bring the overall actor and mention ratios into the (default) target range $[0.75,1.25]$. Compared to the filtered corpus, the distribution is more centred and less polarised, indicating that articles with extreme gender dominance were successfully downsampled to achieve a more balanced overall representation. (Figure 7).
Figure 7: Distribution of gender ratios across articles after corpus-level balancing.
In the final corpus, gender representation is nearly balanced across both mentions and actor counts (Figure 8). This balance is not just a numerical artefact of article exclusion, but reflects a more even distribution across time and actor types. Compared to the original corpus (cf. Figure 1), the trajectories of female- and male-coded representation converge, with both actor and mention lines approaching parity. Importantly, the gender crossing point around 2018 persists, indicating that key corpus dynamics remain intact after balancing.
@@
Figure 8: Percentage of male- and female-coded references over time after all filtering and balancing steps.
Figure 9 shows that the quotation gap is reduced compared to Figure 2. Women now appear in direct speech more frequently than before, while men remain slightly more dominant in indirect speech. The remaining asymmetry may reflect residual stylistic preferences in journalistic practice or shifts in topic domains rather than systematic bias. Overall, this pattern reflects a notable improvement in discursive agency, with women quoted more often in their own words.
@@
Figure 9: Proportion of direct and indirect quotations by gender after full exclusion.
Sentiment remains slightly more negative for women (cf. Figure 10), but the gap remains below the exclusion threshold of 0.3. This persistence may signal that subtle evaluative framing is harder to eliminate via structural balancing. It also suggests that sentiment operates on a different level than grammatical or referential bias, possibly rooted more deeply in lexical associations or broader discourse conventions.
Figure 10: Average sentiment towards she/her and he/him actors over time after full exclusion.
Figure 11 reveals a modest shift in syntactic agency. The proportion of women in subject positions increases slightly, while object roles are more evenly distributed. Although men still dominate in subject roles, the difference is reduced compared to the original corpus. Specifically, the subject-role gap between men and women decreases from approximately 30 to 5 percentage points after filtering. This suggests that our multi-step exclusion process leads to subtle improvements in how grammatical agency is distributed between genders.
@@
Figure 11: Distribution of syntactic roles after full exclusion.
Taken together, these results demonstrate the effectiveness of our multi-stage filtering and balancing strategy in mitigating structural gender asymmetries in the corpus. By comparing before and after exclusion, we observe clear improvements across key dimensions of representation and framing.
Referential parity is achieved: male- and female-coded actors now appear in comparable proportions across both mentions and actor counts. Discursive patterns also show notable shifts. Women are quoted directly more often, and syntactic agency is redistributed more evenly, with a marked reduction in the subject-object gap. These improvements suggest that our approach meaningfully alters how gendered actors are positioned within the narrative structure of the corpus.
However, some subtle forms of bias remain. The sentiment gap persists below the exclusion threshold, indicating that evaluative language is harder to correct through structural rebalancing alone. Such residual asymmetries may reflect deeper, more diffuse biases embedded in lexical or thematic choices rather than in sentence-level grammar.
Overall, the final corpus provides a significantly more equitable foundation for downstream NLP applications and for critical media analysis. By combining large-scale actor-level auditing with targeted corpus interventions, our approach offers a concrete path towards fairer data curation in practice.
@@ 6 Conclusion and Future Work
In this paper, we presented an extended actor-level pipeline for detecting and mitigating gender discrimination in large-scale text corpora. Building on prior work, we introduced new metrics that capture asymmetries in syntactic roles, quote attribution, and sentiment framing. We enhanced the interpretability of the results through structured reports and implemented a two-stage filtering mechanism that enables the construction of gender-balanced corpora.
Our application of the pipeline to the taz2024full corpus demonstrates that gender imbalances in representation and framing are both measurable and correctable to a significant extent. The resulting corpus shows improved balance across multiple linguistic dimensions and serves as a more equitable foundation for downstream tasks such as language model training or critical media studies.
Nevertheless, some forms of discrimination, particularly those tied to sentiment and more implicit discourse structures, persist despite structural balancing. This indicates that not all bias can be addressed through surface-level interventions alone. Future work should therefore explore complementary strategies such as employing context-aware language models for deeper semantic analysis, developing targeted debiasing methods to address persistent framing asymmetries, and integrating intersectional attributes such as race, age, or class. In addition, extending actor categories beyond the gender binary would enable the inclusion of non-binary and gender-diverse identities, allowing for a more comprehensive understanding of representational fairness.
More broadly, we advocate for the integration of discourse-aware methods into standard corpus construction workflows. Understanding how social groups are framed in language is a necessary prerequisite for designing fairer NLP systems, and our pipeline offers a scalable, modular, and linguistically grounded way to do so.
@@
While our approach enables corpus-level balancing based on measurable framing asymmetries, it is not without limitations. First, the exclusion-based strategy necessarily reduces corpus size and diversity, potentially eliminating valuable content alongside discriminatory texts. Second, the method operates on surface-level linguistic signals and cannot fully account for subtler or context-dependent forms of bias, such as irony, framing through omission, or topic selection. Third, the balancing relies on binary gender classification, which excludes non-binary identities and reinforces a gender dichotomy that our conceptual framework otherwise seeks to challenge. Fourth, the analysis is limited to texts that contain clearly identifiable actors and sufficient gender cues, primarily via pronoun resolution. Articles without identifiable pronouns or mentions of coreferent actors are excluded from the discrimination analysis altogether, leading to incomplete coverage. Finally, the impact-aware exclusion method is sensitive to threshold settings and metric selection, which may affect outcomes in ways that are not always transparent. These limitations highlight the need for complementary strategies, such as counterfactual augmentation, contextual bias detection, or narrative-level analysis, to address more complex and nuanced forms of representational inequality.
@@
Our work is grounded in the belief that fairness in NLP requires not only technical interventions but also critical reflection on the social impact of language technologies. By analysing how gendered actors are represented and framed in text, we aim to make structural inequalities visible and address them at the level of data design. At the same time, we recognise that fairness cannot be reduced to numerical balance. The act of filtering texts, however principled, entails normative decisions about which content is considered discriminatory and which is preserved. This introduces risks of over-correction, loss of valuable context, and the potential erasure of complex identities. Furthermore, our reliance on binary gender resolution excludes non-binary, intersex, and gender-nonconforming individuals, reinforcing the very simplifications we seek to critique. We consider this a significant ethical limitation and prioritise extending our methods to support more inclusive representations in future work. Finally, while we aim to mitigate bias in training data, we stress that ethical responsibility must also extend to model architectures, deployment contexts, and the broader socio-technical systems in which NLP tools are embedded.

### v2

@@ 5 Corpus-Balancing of taz2024full
We use the taz2024full corpus Urchs et al. (2025), comprising over 1.8 million articles from the German left-leaning newspaper taz (1980–2024). In the unfiltered corpus, we detect female- and male-coded actors in 1,834,018 articles. Actor frequency peaks in 2004 with 23,580 actors (7,523 female and 16,057 male). In early years, coverage is sparse and dominated by a small number of actors, but from 1988 onwards the corpus broadens significantly.
5.1 Imbalances Before Filtering
Across the unfiltered corpus, men dominate both actor counts and mention frequencies (Figure 1). These asymmetries are reflected not only in absolute representation but also in discursive positioning. Men appear more often in subject roles (cf. Figure 3) and as speakers in direct quotations (cf. Figure 2), while women are more frequently placed in object positions (cf. Figure 3) or paraphrased through indirect quotes (cf. Figure 2).
Figure 1: Percentage of male- and female-coded references over time before filtering. Fluctuations in the early years reflect the small number of available articles.
Figure 2: Distribution of quotation styles by gender before filtering. Early-year fluctuations are attributable to low article counts.
Figure 3: Distribution of syntactic roles by gender before filtering. Early-year fluctuations are attributable to low article counts.
At the article level, gender representation is highly polarised: many texts reference either only male-coded or only female-coded actors (Figure 4). This shows that imbalance is not simply an aggregate effect but is embedded in the composition of individual articles.
@@ 5.1 Imbalances Before Filtering
5.2 Asymmetry Flags
During the first text-level filtering step, we excluded 20 articles using four asymmetry flags: sentiment gap, quote imbalance, subject/object ratio, and naming imbalance (named vs. pronoun mentions). We decided to trigger the document exclusion if two or more flags were detected in a text to prevent over-exclusion.
Each flag compares ratio differences between female- and male-coded actors with +1 Laplace smoothing to avoid division by zero, and fires when the absolute difference exceeds a preset threshold: sentiment gap >0.3>0.3 (difference in average polarity), subject/object ratio difference >0.5>0.5, direct/indirect quote ratio difference >0.5>0.5, and named/pronoun mention ratio difference >0.5>0.5. The cut-off values were chosen to capture asymmetries that go beyond natural stylistic variation and that are likely to affect how actors are framed in discourse. For sentiment, a relatively low threshold of 0.3 was used, since polarity scores are generally close to neutral and even moderate differences can shift evaluative framing. For subject/object roles, quoting, and naming, we required larger divergences of 0.5 in ratio space. These features are structurally more variable across texts, and a stricter cut-off ensures that only sustained imbalances are flagged.
Flag overlap.
Co-occurrence analysis shows that only twenty of the texts exhibit multiple asymmetries simultaneously. Most excluded texts (17,212) are flagged for a single imbalance, primarily subject/object distribution. 564 articles did not trigger any of the four asymmetry flags and were removed in the subsequent corpus-level balancing step to bring the overall actor and mention ratios into the target equilibrium.
Flag frequencies.
Figure 5 shows the share of excluded texts per year by flag type. The subject gap dominates throughout the corpus, consistently accounting for more than 80% of flagged texts across all decades. This stability suggests that structural asymmetries in grammatical agency are a persistent feature of the newspaper’s coverage rather than a phenomenon tied to specific periods. The other three indicators occur more rarely, together contributing less than 10% of exclusions. The quote gap shows the most variation over time: it reaches values of up to 5-6% of excluded texts in the 1990s and early 2000s, but remains lower and more stable after 2010. These spikes may reflect topic-specific reporting practices in those decades, such as an emphasis on political debates or international conflicts where male actors dominated as attributed speakers, while female actors were more often paraphrased. The naming gap occurs at low levels (1-2%) without a clear temporal trend, while the sentiment gap is negligible throughout, with only a slight increase visible after 2010. Overall, no systematic long-term trends are observable beyond the persistent dominance of subject-role asymmetries and the temporary spikes in quoting imbalances around the turn of the millennium.
Figure 5: Proportion of excluded texts per year by flag type. Subject-role asymmetry dominates, while naming, quoting, and sentiment gaps occur less frequently.
Qualitative examples.
To illustrate these asymmetries, we include examples from the excluded set:
  • Subject-object gap: In a football report on Eintracht Frankfurt, all named actors are male and consistently appear as grammatical subjects: “Horst Ehrmantraut […] gelang es, mit geringen finanziellen Mitteln den Aufstieg zu realisieren”, “Rolf Heller […] regiert heute auf dem Präsidentenstuhl”, “Weber […] hat nach langem Pokern einen neuen Vierjahresvertrag unterschrieben”.
    Female-coded actors are entirely absent from the text, reinforcing an imbalance where men hold agency in the discourse while women do not appear as subjects at all.
  • Quote imbalance: In a political portrait of Peter-Michael Diestel, male actors are repeatedly given direct speech: “Alle, alle, waren da und wollten mich haben […]”, “Ich bin strunzbieder. Ich bin ein Konservativer. Ich stehe zum CDU-Programm.”, “Vor Schröder hätte er Schiß gehabt.”
    Female-coded actors, by contrast, are only mentioned collectively (e.g. “Eppelmann, Heitmann und andere […]”) and paraphrased without direct quotations.
  • Naming gap: In a film review, the female protagonist is repeatedly introduced by name: “Deniz, die 21-jährige Heldin in Arslans Film, geht an diesem nicht enden wollenden Sommertag Rohmer-Filme synchronisieren., "Wie Deniz an diesem Tag ihre Wäsche zur Mutter bringt, die Schwester trifft, ihren Freund verlässt […]”
    Male figures in the same text, such as her boyfriend or the director Thomas Arslan, are mentioned once and then largely referred to with pronouns.
  • Sentiment gap: In a letter to the editor, female actors are explicitly evaluated in negative terms, for example: “[…] wie ist es möglich, dass die Autorin ohne Kommentar oder Richtigstellung wahrheitswidrig schreiben kann […]” and dismissively mocked: “Sie sind herzlich eingeladen, für ihre hehren Werte mit einer Menschenkette an der Front […] zu demonstrieren.”
    By contrast, male commentators in the same article (e.g. Hartmut Rosa, Peter Bethke, Gerhard Harms) are described neutrally or respectfully.
5.3 Global Equilibrium
In a second step, we applied corpus-level balancing, excluding an additional 17,816 articles to bring the overall actor and mention ratios into the target range $[0.75,1.25]$. This interval means that one gender may occur up to 25% more frequently than the other: for example, a ratio of 1.25 indicates that female actors or mentions outnumber male ones by 25%, while a ratio of 0.75 indicates the reverse. Both directions are treated symmetrically, ensuring that neither male nor female dominance persists beyond this margin. The choice of range enforces approximate parity without creating an artificial 1:1 distribution, while retaining authentic temporal dynamics.
Figure 6 shows the resulting distribution of gender ratios across all articles. Compared to the unfiltered corpus (cf. Figure 4), the distribution is more centred and less polarised. Articles with exclusively male-coded or exclusively female-coded actors, which previously created sharp spikes at 0% and 100%, have been reduced. Instead, more texts now fall into the mid-range, where both genders are present. This demonstrates that global balancing successfully decreased the extreme ends of the distribution while preserving variation in the middle range.
Figure 6: Distribution of gender ratios across articles after corpus-level balancing. The x-axis shows the percentage of she/her references (mentions on the left, actors on the right). The y-axis shows the proportion of texts. Peaks at 0% and 100% are strongly reduced after balancing, indicating that one-gender-only articles were downsampled.
Excluded texts.
The equilibrium step removed texts that were maximally polarised in their gender representation. Across all 17,796 excluded articles, we detect 35,995 male-coded actors and 190,192 male-coded mentions, but no female-coded actors or mentions. In other words, every article excluded by this step contained only men. Such one-sided texts were frequent enough to skew the corpus-level balance if left untouched, producing systematic over-representation of male-coded actors.
Temporal distribution.
Exclusions occur across the entire corpus history (Figure 7), but their frequency closely tracks overall article production. In the late 1980s and early 1990s, very few articles are excluded, reflecting the limited size of the corpus at that stage. From the mid-1990s onwards the number of exclusions rises steadily, stabilising at around 400–500 per year. Between 2005 and 2015 exclusions remain consistently high, often exceeding 500 texts annually, with a clear peak around 2015 (over 650). After 2018, the numbers decline again, falling below 300 in the most recent years. This pattern indicates that exclusions are not confined to the early, sparse period, but accompany phases of high article production and decline in line with overall corpus dynamics. Furthermore, we could detect no temporal trend in discrimination.
Figure 7: Number of excluded texts per year after all exclusion steps (text-level filtering and corpus-level balancing). Exclusions scale with article production and are distributed across the corpus history.
5.4 Results After Filtering and Balancing
The final corpus exhibits near parity across mentions and actor counts (Figure 8). Importantly, structural dynamics such as the crossing point around 2018 remain intact, indicating that balancing improves representation without erasing genuine historical patterns. Compared to the unfiltered corpus (cf. Figure 1), the trajectories of male- and female-coded actors now run in parallel, showing that referential balance has been restored across time.
@@ 5.4 Results After Filtering and Balancing
Figure 8: Percentage of male- and female-coded references over time after all filtering and balancing steps. The trajectories of mentions and actors converge towards parity while retaining the natural crossing point around 2018.
Figure 9 shows the distribution of quotation styles. After balancing, women appear more often in direct speech than before, reducing the quote imbalance observed in Figure 2. Men still receive slightly more indirect quotations, but the gap is narrower, suggesting that women’s discursive agency is more strongly represented in the final corpus.
@@
Figure 9: Proportion of direct and indirect quotations by gender after full exclusion. The gap is reduced compared to the unfiltered corpus, with women more frequently quoted directly.
Finally, Figure 10 illustrates the distribution of syntactic roles. Men still occur more frequently in subject positions, but the difference is markedly reduced compared to the unfiltered corpus (cf. Figure 3). The gap decreases from around 30 percentage points to roughly 5, showing that grammatical agency is now distributed more evenly across genders. This represents one of the strongest structural improvements achieved by the balancing process.
@@
Figure 10: Distribution of syntactic roles after full exclusion. The subject–object gap between men and women is markedly reduced compared to the unfiltered corpus.
Taken together, these figures demonstrate that the corpus is not only numerically balanced but also structurally improved. Referential parity is achieved, women are quoted more often in their own words and syntactic agency is redistributed more evenly. The balancing process thus mitigates multiple dimensions of gender inequality while preserving historically meaningful variation.
@@ 6 Conclusion and Future Work
We presented an extended actor-level pipeline for detecting and mitigating gender discrimination in large-scale text corpora. Beyond prior work, we introduced metrics for syntactic roles, quotation, and sentiment, structured reports for interpretability, and a two-stage filtering process for building more balanced corpora.
Applied to the taz2024full corpus, our approach shows that gender imbalances in representation and framing are both measurable and correctable. The resulting corpus is more balanced across multiple linguistic dimensions and provides a stronger foundation for corpus-based analysis and fairer NLP practices.
Yet some asymmetries, particularly in implicit discourse structures, persist. Future work should address these through context-aware models, targeted debiasing strategies, and intersectional extensions that include race, age, and class. Expanding actor categories beyond the gender binary will further support inclusive analysis. More broadly, we argue that discourse-aware methods should become part of corpus construction workflows, as understanding how groups are framed is essential for designing fairer NLP systems.
@@
While our approach enables corpus-level balancing based on measurable framing asymmetries, it has limitations. The exclusion strategy reduces corpus size and may remove valuable content alongside biased texts. It also relies on surface-level linguistic signals and cannot capture subtler biases such as irony, omission, or topic choice. Furthermore, the method enforces a binary gender classification, excluding non-binary identities, and it applies only to texts with identifiable actors and gender cues, leaving some material outside the analysis.
@@
Our work is grounded in the belief that fairness in NLP requires not only technical interventions but also critical reflection on the social impact of language technologies. By analysing how gendered actors are represented and framed in text, we make structural inequalities visible and address them at the level of data design. Yet fairness cannot be reduced to numerical balance: filtering texts entails normative choices about which content is deemed discriminatory, with risks of over-correction and loss of context. Our reliance on binary gender resolution further excludes non-binary and gender-nonconforming individuals, reinforcing the very simplifications we seek to critique. We consider this a significant ethical limitation and aim to extend our methods to more inclusive representations. Finally, while we mitigate discrimination in training data, responsibility also lies in model architectures, deployment contexts, and the socio-technical systems in which NLP tools operate.